A direct compressive force of 240,000 N is applied to a solid piece of material of rectangular cross-section, 8 cm by 5 cm.
1
Expert's answer
2020-04-08T10:25:56-0400

Given that,

A direct compressive force of 240,000 N is applied to a solid piece of material of rectangular cross-section, 8 cm by 5 cm.


Therefore the total cross sectional area = "8cm\times5cm=40cm^2"


Pressure acting on the solid piece of material by the compressive force is given by,


"P=\frac{F}{A}=\frac{240000}{40 \times 10^{-4}}"

"=6\times10^{7}Nm^{-2}"